Some students planned for yoga Camp. The total budget was 240 Duet. Some reason 4 students failed to attends the camp and thus the cost of food for each member increased by Rs 5 . How many students attend the yoga camp and how much each student .pay for food? what value of students are depicted in the question by their activity?

Grade:10

1 Answers

Hadiya
6 years ago
Let X be the number of students who plan to attend the camp.Let Y be the amount each student have to pay.XY=240Y=240/XSince 4 Students fail to attend,(X-4)(Y+5)=240(given)(X-4)(240/X. +5)=240Solving this we get,5X^2-20X-960=0Solving this quadratic equation we get,X=16Y=15So the number of students who attended the camp =16-4=12.Amount each student who attended paid for food=15+5=20Regards,Hadiya